OpenStudy (anonymous):

Please help: How can you write the expression w/a rationalized denominator. 5-√2 / √3

OpenStudy (solomonzelman):

\[\frac{5- \sqrt{2}}{\sqrt{3}}\]top and bottom times square root of 3.

OpenStudy (solomonzelman):

\[\frac{(5-\sqrt{2}) \times \sqrt{3}}{\sqrt{3} \times \sqrt{3}}=\frac{5\sqrt{3}-\sqrt{6}}{3}\]